# Optical systems with translational invariance

### v2

@@ 3.1 Application of Fermat’s Principle
$\alpha_{c1}$ $=$ $-\frac{3}{2}\left(n_{0}\frac{\sin\theta_{0}}{s_{0}^{2}}\Theta_{0}^{x}-n_{1}\frac{\sin\theta_{1}}{s_{1}^{2}}\Theta_{1}^{x}\right),\quad\alpha_{c2}=-\frac{1}{2}\left(\frac{n_{0}\sin\theta_{0}}{s_{0}^{2}}-\frac{n_{1}\beta^{2}\sin\theta_{1}}{s_{1}^{2}}\right),$ (60)
$\alpha_{a1}$ $=$ $\frac{n_{0}}{s_{0}}\Theta_{0}^{x}+\frac{n_{1}}{s_{1}}\Theta_{1}^{x},\quad\alpha_{a2}=\frac{n_{0}}{s_{0}}-\frac{n_{1}}{s_{1}}(1+2\beta),\quad\alpha_{d}=n_{0}\sin\theta_{0}-n_{1}\sin\theta_{1}$ (61)
@@
$\delta_{1x}$ $=$ $\frac{s_{1}}{n_{1}}\left[(\alpha_{e1}x^{\prime 2}+\alpha_{e3}y^{\prime 2})x^{\prime}+(\alpha_{c1}x^{\prime 2}+\alpha_{c2}y^{\prime 2})+\alpha_{a1}x^{\prime}+\alpha_{d}\right],$ (63)
@@ 3.2 Evaluation of aberration coefficients
The aberration functions in Eqs. (58)–(61) can be evaluated as follows. For the terms representing coma we replace each cosine by one and each sine by its angle. Using Eq. (43), we have
@@
Again approximating cosines by unity and sines by their angles, using Eqs. (67) and the leading order approximation of the Snell’s law
@@
$\alpha_{e2}$ $=$ $-\frac{n_{0}}{2s_{0}^{3}}\left(1-\frac{s_{0}^{2}}{s_{1}^{2}}\,\frac{1+4\beta+6\beta^{2}+4\beta^{3}}{\beta}\right)\quad{\rm with}\quad\beta=\frac{n_{0}s_{1}}{n_{1}s_{0}},$ (71)
@@
$\alpha_{a1}$ $=$ $-\frac{n_{0}\theta_{0}^{2}}{s_{0}}\left(1+\frac{n_{0}s_{0}}{n_{1}s_{1}}\right)-\frac{n_{0}\theta_{0}^{2}}{2R}\left(1-\frac{n_{0}}{n_{1}}\right),$ (74)
$\alpha_{a2}$ $=$ $-\frac{n_{0}}{s_{0}}\left(1+\frac{s_{0}}{n_{0}}\frac{n_{1}}{s_{1}}\right)-\frac{n_{0}\theta_{0}^{2}}{s_{0}}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right).$ (75)
In consideration of $\alpha_{d}$ the first order approximation of the Snell’s law in Eq. (69) is not sufficient. In order to preserve the precise direction from which the chief ray approaches the stop (and the object and image planes), it is necessary to expand sines and tangents to higher orders [14]. From Eqs. (40)-(41) we have $\tan\theta_{0}=x_{0}/z_{0}$ and $\tan\theta_{1}=x_{1}/z_{1}$. To third order
@@
$\alpha$ Refracting surface of Eq. (38) Reflecting surface of Eq. (80)
@@
$\alpha_{e2}$  $-\frac{n_{0}}{2s_{0}^{3}}\left(1-\frac{s_{0}^{2}}{s_{1}^{2}}\,\frac{1+4\beta+6\beta^{2}+4\beta^{3}}{\beta}\right)$  $-\frac{n_{0}}{2s_{0}^{3}}\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{\mu^{3}}$
@@
$\alpha_{a2}$  $-\frac{n_{1}-n_{0}}{R}-\frac{n_{0}\theta_{0}^{2}}{s_{0}}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)$  $-\frac{n_{0}}{s_{0}}\frac{\mu-1}{\mu}$
@@ 3.3 Aberrations for a mirror surface
$\Theta_{i}^{x}$ $=$ $\left(1-\frac{s_{i}}{R}\right)-\frac{\theta_{0}^{2}}{2}\left(2-\frac{s_{i}}{R}\right),\qquad\quad\Theta_{i}^{xy}=-\left(1-\frac{s_{i}}{R}\right)+\frac{\theta_{0}^{2}}{2}\left(6-\frac{s_{i}}{R}\right),$ (84)
@@
For the mirror surface the function $\beta$ in Eq. (51) reduces to the transverse (lateral) magnification
@@
using Eqs. (41) and (81), the analog of Eq. (49) writes as
@@
Hence in Eq. (56) $1+\beta$ is replaced by $1-\mu$ and the expressions for $\alpha_{a2}$ and $\alpha_{e2}$ read as
$\alpha_{e2}=-\frac{n_{0}}{2s_{0}^{3}}+\frac{n_{1}}{2s_{1}^{3}}\left[(1-\mu)^{4}-\mu^{4}\right],\qquad\alpha_{a2}=\frac{n_{0}}{s_{0}}-\frac{n_{1}}{s_{1}}(1-2\mu).$ (89)
@@
and derive the expressions for $\alpha_{i}$ shown in the third column of Table 1. The distortion coefficient $\alpha_{d}$ is vanishing same as for a mirror of revolution [14]. However, in the case of cylindrical mirror the second coma is also vanishing: $\alpha_{c2}=0$. Hence, in this case only two additional aberrations left with respect to axially symmetric surface, and these two new degrees of freedom modify the spherical aberration.
(Possibilities of compensation this modified spherical aberration for the systems of cylindrical mirrors are investigating in the following).
@@ 3.4 Aberrations for displaced stop
Figure 3: Portion of Fig. 1 with aperture stop displaced from surface by a distance $W$ along the $Z$ axis. The entrance pupil is at the stop. The chief ray makes angle $\psi$ with the $Z$ axis and intersects the surface at height $L$.
Consider a single optical surface with the aperture stop displaced from it by a distance $W$, as shown in Fig. 3. The chief ray intersects the surface at height $L_{x}$. Due to the lack of rotational symmetry this is not enough to make the replacement of $x^{\prime}\to x^{\prime}+L_{x}$ in the wave aberration as usually [14]. Now the complete change to $\Phi(x^{\prime}+L_{x},y^{\prime}+L_{y})$ should be considered. However, taking into account that the ray behavior in the $XZ$ plane is the same in the discussed cylindrical OS as in the correspondent rotationally symmetric OS (with the same $XZ$ profile), we consider the $\Phi(x^{\prime}+L_{x},y^{\prime})$ part first to simplify further comparisons to the classical results.
@@ 3.4.1 $L_{y}=0$ part of the aberration function
Making the replacement of $x^{\prime}\to x^{\prime}+L_{x}$ in Eq. (57) and collecting terms in various powers of $x^{\prime}$ and $y^{\prime}$, we get
$\Phi^{(4)}(x^{\prime}+L_{x},y^{\prime})$ $=$ $A_{01}x^{\prime}+A_{11}x^{\prime 2}+A_{12}y^{\prime 2}+A_{21}x^{\prime 3}+A_{22}x^{\prime}y^{\prime 2}$ (92)
$+$ $A_{31}x^{\prime 4}+A_{32}y^{\prime 4}+A_{33}x^{\prime 2}y^{\prime 2}+C_{0},$
@@
$A_{11}$ $=$ $\frac{\alpha_{a1}}{2}+\alpha_{c1}L_{x}+\frac{3\alpha_{e1}}{2}L_{x}^{2},\qquad\quad A_{12}=\frac{\alpha_{a2}}{2}+\alpha_{c2}L_{x}+\frac{\alpha_{e3}}{2}L_{x}^{2},$ (94)
$A_{21}$ $=$ $\frac{\alpha_{c1}}{3}+\alpha_{e1}L_{x},\qquad\quad A_{22}=\alpha_{c2}+\alpha_{e3}L_{x},$ (95)
$A_{31}$ $=$ $\frac{\alpha_{e1}}{4},\qquad\quad A_{32}=\frac{\alpha_{e2}}{4},\qquad\quad A_{33}=\frac{\alpha_{e3}}{2}.$ (96)
@@
$2\theta_{0}^{2}+6\theta_{0}L_{x}\Gamma_{R}+3L_{x}^{2}\Gamma_{R}^{2}$ $=$ $(W\psi)^{2}\left[3\left(\frac{1}{W}\pm\frac{1}{R}\right)^{2}-\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\right].$ (103)
@@
From the results for $\alpha_{i}$ in Table 1 and Eqs. (101)–(103), we derive the expressions for the coefficients $A_{ij}$ shown in Table 2, where the upper and lower signs correspond to the stop located in front of (see Fig. 3) and behind the surface, respectively.
Table 2: Aberration coefficients for cylindrical surface with displaced stop and $L_{y}=0$
$A_{ij}$ Refracting surface of Eq. (38) Reflecting surface of Eq. (80)
$A_{31}$  $-\frac{n_{0}}{8}\left(\frac{\kappa}{R^{3}}\frac{n_{1}-n_{0}}{n_{0}}+b+\Gamma\Gamma_{R}^{2}\right)$  $-\frac{n_{0}}{4R^{3}}\left[\kappa+\left(\frac{\mu+1}{\mu-1}\right)^{2}\right]-\frac{b}{8}$
$A_{32}$  $-\frac{n_{0}}{8s_{0}^{3}}\left[1-\frac{s_{0}^{2}}{s_{1}^{2}}\,\frac{(1+\beta)^{4}-\beta^{4}}{\beta}\right]$  $-\frac{n_{0}}{8s_{0}^{3}}\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{\mu^{3}}$
$A_{33}$  $-\frac{n_{0}}{4s_{0}^{2}}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\Gamma_{R}$  $-\frac{n_{0}}{2s_{0}^{2}R}\,\frac{\mu+1}{\mu-1}$
$A_{21}$  $-\frac{n_{0}}{2}(W\psi)\left[\pm\frac{\kappa}{R^{3}}\frac{n_{1}-n_{0}}{n_{0}}\pm b+\left(\frac{1}{W}\pm\frac{1}{R}\right)\Gamma\Gamma_{R}\right]$  $-\frac{n_{0}}{R^{2}}(W\psi)\left[\left(\frac{1}{W}\mp\frac{1}{R}\right)\frac{\mu+1}{\mu-1}\pm\frac{\kappa}{R}\right]\mp\frac{b}{2}(W\psi)$
$A_{22}$  $-\frac{n_{0}}{2s_{0}^{2}}(W\psi)\left(\frac{1}{W}\pm\frac{1}{R}\right)\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)$  $\mp\frac{n_{0}}{2s_{0}^{2}R}\left(W\psi\right)\frac{\mu+1}{\mu-1}$
$A_{11}$  $-\frac{n_{0}}{4}(W\psi)^{2}\left[3\left(\frac{1}{W}\pm\frac{1}{R}\right)^{2}-\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\right]\Gamma$  $-\frac{n_{0}}{2R}(W\psi)^{2}\left[3\left(\frac{1}{W}\mp\frac{1}{R}\right)^{2}-2\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}+\frac{3\kappa}{R^{2}}\right]$
 $-\frac{n_{0}}{4}(W\psi)^{2}\left[\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\frac{n_{1}-n_{0}}{n_{1}R}+\frac{3\kappa}{R^{3}}\frac{n_{1}-n_{0}}{n_{0}}+\frac{3b}{n_{0}}\right]$  $-\frac{3b}{4}(W\psi)^{2}$
$A_{12}$  $-\frac{n_{1}-n_{0}}{2R}-\frac{n_{0}}{2s_{0}}(W\psi)^{2}\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)$  $-\frac{n_{0}}{2s_{0}}\frac{\mu-1}{\mu}-\frac{n_{0}}{2s_{0}^{2}R}\left(W\psi\right)^{2}\frac{\mu+1}{\mu-1}$
 $-\frac{n_{0}}{2s_{0}^{2}}(W\psi)^{2}\left(\frac{1}{2R}\pm\frac{1}{W}-\frac{1}{2s_{0}}\right)\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)$
$A_{01}$  $-\frac{n_{0}}{2}(W\psi)^{3}\left\{\left(\frac{1}{W}\pm\frac{1}{R}\right)\left(\frac{1}{R}\pm\frac{2}{W}-\frac{1}{s_{0}}\right)\Gamma\right.$  $\mp\frac{n_{0}}{2R}(W\psi)^{3}\left[3\left(\frac{1}{W}\mp\frac{1}{R}\right)^{2}-\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\right]$
 $+\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\left[\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\pm\frac{n_{1}-n_{0}}{n_{1}R}\right]$  $\mp\frac{n_{0}}{2R^{3}}(W\psi)^{3}\left[\kappa-\left(\frac{\mu+1}{\mu-1}\right)^{2}\right]\mp\frac{b}{2}(W\psi)^{3}$

### v3

@@ 3.1 Application of Fermat’s Principle
$\alpha_{c1}$ $=$ $-\frac{3}{2}\left(n_{0}\frac{\sin\theta_{0}}{s_{0}^{2}}\Theta_{0}^{x}-n_{1}\frac{\sin\theta_{1}}{s_{1}^{2}}\Theta_{1}^{x}\right),$ (58)
$\alpha_{c2}$ $=$ $-\frac{1}{2}\left(\frac{n_{0}\sin\theta_{0}}{s_{0}^{2}}-\frac{n_{1}\beta^{2}\sin\theta_{1}}{s_{1}^{2}}\right),$ (59)
$\alpha_{a1}$ $=$ $\frac{n_{0}}{s_{0}}\Theta_{0}^{x}+\frac{n_{1}}{s_{1}}\Theta_{1}^{x},$ (60)
$\alpha_{d}$ $=$ $n_{0}\sin\theta_{0}-n_{1}\sin\theta_{1}$ (61)
@@
$\delta_{1x}=\frac{s_{1}}{n_{1}}\left[(\alpha_{e1}x^{\prime 2}+\alpha_{e3}y^{\prime 2})x^{\prime}+(\alpha_{c1}x^{\prime 2}+\alpha_{c2}y^{\prime 2})+\alpha_{a1}x^{\prime}+\alpha_{d}\right],$
@@ 3.2 Evaluation of aberration coefficients
The aberration functions in Eqs. (56)–(61) can be evaluated as follows. For the terms representing coma we replace each cosine by one and each sine by its angle. Using (43), we have
@@
Again approximating cosines by unity and sines by their angles, using (66) and the leading order approximation of the Snell’s law
@@
$\alpha_{a1}$ $=$ $-\frac{n_{0}\theta_{0}^{2}}{s_{0}}\left(1+\frac{n_{0}s_{0}}{n_{1}s_{1}}\right)-\frac{n_{0}\theta_{0}^{2}}{2R}\left(1-\frac{n_{0}}{n_{1}}\right).$ (72)
In consideration of $\alpha_{d}$ the first order approximation of the Snell’s law in (68) is not sufficient. In order to preserve the precise direction from which the chief ray approaches the stop (and the object and image planes), it is necessary to expand sines and tangents to higher orders [14]. From Eqs. (40)–(41) we have $\tan\theta_{0}=x_{0}/z_{0}$ and $\tan\theta_{1}=x_{1}/z_{1}$. To third order
@@
$\!\alpha$ Refracting surface of (38) Reflecting surface, (77)
@@ 3.3 Aberrations for a mirror surface
$\Theta_{i}^{x}=\left(1-\frac{s_{i}}{R}\right)-\frac{\theta_{0}^{2}}{2}\left(2-\frac{s_{i}}{R}\right),\leavevmode\nobreak\ \Theta_{i}^{xy}=-\left(1-\frac{s_{i}}{R}\right)+\frac{\theta_{0}^{2}}{2}\left(6-\frac{s_{i}}{R}\right)\!,$ (81)
@@
For the mirror surface the function $\beta$ in (50) reduces to the transverse (lateral) magnification
@@
using Eqs. (41) and (78), the analog of (48) writes as
@@
and derive the expressions for $\alpha_{i}$ shown in the third column of Table 1. The distortion coefficient $\alpha_{d}$ is vanishing same as for a mirror of revolution [14]. However, in the case of cylindrical mirror the second coma is also vanishing: $\alpha_{c2}=0$. Hence, in this case only two additional aberrations left with respect to the case of axially symmetric surface, and these two new degrees of freedom modify the spherical aberration. Possibilities for compensation of the modified spherical aberration in the systems of cylindrical mirrors are discussed in section 4.
@@ 3.4 Aberrations for displaced stop
Figure 4: Upper [Lower]: portion of Fig. 2 [3] with aperture stop displaced from surface by a distance $W$ along the $Z$ axis. The entrance pupil is at the stop. The chief ray makes angle $\psi$ [$\chi\approx\eta_{0}$] with the $Z$ axis and intersects the surface at height $L_{x}$ [$L_{y}$].
Consider a single optical surface and the aperture stop displaced from it by a distance $W$, as shown in Fig. 4. The chief ray, which is now $P_{0}P^{\prime}$, is directed towards the pupil’s center and intersects the surface at heights $L_{x}$ and $L_{y}$ in the $XZ$ and $YZ$ planes, respectively, and the wave aberration writes as $\Phi(x^{\prime}+L_{x},y^{\prime}+L_{y})$. However, taking into account the fact that the ray behavior in the $XZ$ plane is the same in the discussed cylindrical OS as in the corresponding rotationally symmetric OS with the same $XZ$ profile, we consider the $\Phi(x^{\prime}+L_{x},y^{\prime})$ part first to simplify further comparisons to the classical results [14].
@@ 3.4.1 $L_{y}=0$ part of the aberration function
Making the replacement of $x^{\prime}\to x^{\prime}+L_{x}$ in (55) and collecting terms in various powers of $x^{\prime}$ and $y^{\prime}$, we get
$\Phi^{(4)}(x^{\prime}+L_{x},y^{\prime})$ $=A_{01}x^{\prime}+A_{11}x^{\prime 2}+A_{12}y^{\prime 2}+A_{21}x^{\prime 3}+A_{22}x^{\prime}y^{\prime 2}$
$+A_{31}x^{\prime 4}+A_{32}y^{\prime 4}+A_{33}x^{\prime 2}y^{\prime 2}+C_{0},$ (88)
@@
$A_{11}$ $=\frac{\alpha_{a1}}{2}+\alpha_{c1}L_{x}+\frac{3\alpha_{e1}}{2}L_{x}^{2},\quad A_{12}=\alpha_{c2}L_{x}+\frac{\alpha_{e3}}{2}L_{x}^{2},$ (90)
$A_{21}$ $=\frac{\alpha_{c1}}{3}+\alpha_{e1}L_{x},\quad A_{22}=\alpha_{c2}+\alpha_{e3}L_{x},$ (91)
$A_{31}$ $=\frac{\alpha_{e1}}{4},\quad A_{32}=0,\quad A_{33}=\frac{\alpha_{e3}}{2}.$ (92)
@@
$2\theta_{0}^{2}+6\theta_{0}L_{x}\Gamma_{R}+3L_{x}^{2}\Gamma_{R}^{2}$ $=(W\psi)^{2}$
×$\times\left[3\left(\frac{1}{W}\pm\frac{1}{R}\right)^{2}-\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\right].$ (99)
@@
From the results for $\alpha_{i}$ in Table 1 and Eqs. (97)–(99), we derive the expressions for the coefficients $A_{ij}$ shown in Tables 2 and 3, where the upper and lower signs correspond to the stop located in front of (see Fig. 4) and behind the surface, respectively.
Table 2: Aberration coefficients for the refracting cylindrical surface of (38) with displaced stop and $L_{y}=0$
$A_{31}$  $-\frac{n_{0}}{8}\left(\frac{\kappa}{R^{3}}\frac{n_{1}-n_{0}}{n_{0}}+\frac{b}{n_{0}}+\Gamma\Gamma_{R}^{2}\right)$
$A_{33}$  $-\frac{n_{0}}{4s_{0}^{2}}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\Gamma_{R}$
$A_{21}$  $-\frac{n_{0}}{2}(W\psi)\left[\pm\frac{\kappa}{R^{3}}\frac{n_{1}-n_{0}}{n_{0}}\pm\frac{b}{n_{0}}+\left(\frac{1}{W}\pm\frac{1}{R}\right)\Gamma\Gamma_{R}\right]$
$A_{22}$  $-\frac{n_{0}}{2s_{0}^{2}}(W\psi)\left(\frac{1}{W}\pm\frac{1}{R}\right)\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)$
$A_{11}$  $-\frac{n_{0}}{4}(W\psi)^{2}\left[3\left(\frac{1}{W}\pm\frac{1}{R}\right)^{2}-\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\right]\Gamma$
 $-\frac{n_{0}}{4}(W\psi)^{2}\left[\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\frac{n_{1}-n_{0}}{n_{1}R}+\frac{3\kappa}{R^{3}}\frac{n_{1}-n_{0}}{n_{0}}+\frac{3b}{n_{0}}\right]$
$A_{12}$  $-\frac{n_{0}}{2s_{0}^{2}}(W\psi)^{2}\left(\frac{1}{2R}\pm\frac{1}{W}-\frac{1}{2s_{0}}\right)\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)$
$A_{01}$  $-\frac{n_{0}}{2}(W\psi)^{3}\left\{\left(\frac{1}{W}\pm\frac{1}{R}\right)\left(\frac{1}{R}\pm\frac{2}{W}-\frac{1}{s_{0}}\right)\Gamma\right.$
 $+\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\left[\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\pm\frac{n_{1}-n_{0}}{n_{1}R}\right]$
@@
Table 3: Aberration coefficients for the reflecting cylindrical surface of (77) with displaced stop and $L_{y}=0$
$A_{31}$  $-\frac{n_{0}}{4R^{3}}\left[\kappa+\left(\frac{\mu+1}{\mu-1}\right)^{2}\right]-\frac{b}{8}$
$A_{33}$  $-\frac{n_{0}}{2s_{0}^{2}R}\,\frac{\mu+1}{\mu-1}$
$A_{21}$  $-\frac{n_{0}}{R^{2}}(W\psi)\left[\left(\frac{1}{W}\mp\frac{1}{R}\right)\frac{\mu+1}{\mu-1}\pm\frac{\kappa}{R}\right]\mp\frac{b}{2}(W\psi)$
$A_{22}$  $\mp\frac{n_{0}}{2s_{0}^{2}R}\left(W\psi\right)\frac{\mu+1}{\mu-1}$
$A_{11}$  $-\frac{n_{0}}{2R}(W\psi)^{2}\left[3\left(\frac{1}{W}\mp\frac{1}{R}\right)^{2}-2\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}+\frac{3\kappa}{R^{2}}\right]$
 $-\frac{3b}{4}(W\psi)^{2}$
$A_{12}$  $-\frac{n_{0}}{2s_{0}^{2}R}\left(W\psi\right)^{2}\frac{\mu+1}{\mu-1}$
$A_{01}$  $\mp\frac{n_{0}}{2R}(W\psi)^{3}\left[3\left(\frac{1}{W}\mp\frac{1}{R}\right)^{2}-\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)^{2}\right]$
 $\mp\frac{n_{0}}{2R^{3}}(W\psi)^{3}\left[\kappa-\left(\frac{\mu+1}{\mu-1}\right)^{2}\right]\mp\frac{b}{2}(W\psi)^{3}$